### v3

# PROMA: Projected Microbatch Accumulation for Reference-Free Proximal Policy Updates

Abstract
This note introduces Projected Microbatch Accumulation (PROMA), a proximal policy method that modifies gradient accumulation across microbatches rather than relying on likelihood ratios relative to a reference policy. During accumulation, PROMA projects the partially accumulated gradient to be orthogonal to the sequence-wise gradients of the current microbatch. This projection is applied layer-wise during the backward pass, enabling efficient implementation. A within-microbatch variant (Intra-PROMA) acts independently across microbatches. Empirically, PROMA achieves proximal updates without entropy collapse while providing tighter local KL control than GRPO.
1 Background
Proximal Policy Optimization (PPO) [9] and its variants, including GRPO [10], stabilize reinforcement learning by limiting the magnitude of policy updates between optimization steps. This is typically achieved by optimizing a surrogate objective that discourages large deviations from a reference policy $\pi_{\text{old}}$ through clipping of the policy likelihood ratio. GRPO and related methods [10, 14] combine PPO-style updates with a group-relative advantage estimator and have become the dominant approach for reinforcement learning–based fine-tuning of large language models. CISPO [3] is another proximal policy variant that clips the importance-sampling ratio in the loss while ensuring that all tokens contribute to the gradient computation. More broadly, proximal policy methods can be viewed as tractable approximations to the natural policy gradient [7], whose exact computation is infeasible at scale.
2 Projected Microbatch Accumulation (PROMA)
PROMA modifies the optimization process during the backward pass. For each microbatch, the sequence-wise log-probability gradients (mcb_seq_grads in Algorithm 2) are computed as part of evaluating the policy gradient, denoted mcb_policy_grad. Before gradients are accumulated, the running gradient t.grad of a parameter tensor t is updated by projecting out components that lie in the span of the sequence-wise log-probability gradients. This projection is applied layer-wise and acts directly on the accumulated gradient.
1def accumulate_PROMA(t, mcb_policy_grad, mcb_seq_grads)
2 """
3 k: number of sequences in microbatch
4 t: (d,) tensor of model weights in the layer
5 mcb_policy_grad: (d,) vanilla policy gradient w.r.t. tensor t for the microbatch
6 mcb_seq_grads: (d, k), sequence-wise gradients of log_prob.sum() w.r.t tensor t
7 """
8 complemented_acc_grad = project_to_complement(t.grad, mcb_seq_grads)
9 t.grad = complemented_acc_grad + mcb_policy_grad
10
11def project_to_complement(acc_vec, vecs):
12 Q, _ = torch.linalg.qr(vecs, mode="reduced") # Q: (d, k)
13 return acc_vec - Q @ (Q.T @ acc_vec)
Figure 1: accumulate_PROMA is called during the backward pass of the policy gradient loss and modifies the gradient accumulation. Before adding the microbatch gradient, the partially accumulated gradient is projected to be orthogonal to the gradient of each sequence in the current microbatch (mcb_seq_grads).
The motivation for PROMA can be put simply: $\nabla\log\pi(s)$ for a sequence $s$ represents the parameter direction where the sequence’s probability changes the fastest. The parameter update $d\theta$ represented by t.grad in Algorithm 2 incurs a KL divergence that grows with the squared overlaps $\mathbb{E}_{s}[(d\theta\;\cdot\;\nabla_{\theta}\log\pi(s))^{2}]$. In Algorithm 2, each column of mcb_seq_grads represents such a sample $\nabla_{\theta}\log\pi(s))^{2}$. PROMA aims to avoid this contribution to the KL divergence by projecting away these components of the gradient update.
Efficiency
To exactly project away the $k$ sequences in the microbatch takes $\approx 2k^{2}d$ FLOPS where $d$ is the number of model parameters. If the microbatch size is large then this can be applied group-wise so that $k\sim 8$. The cost $2k^{2}d$ of PROMA is then negligible compared to the $T\times d$ operations required to form the gradient, where $T$ is the number of tokens. Another option is to apply an approximate projection as in Algorithm 4 which takes $O(kd)$ operations.
3 An intra-microbatch variant (Intra-PROMA)
While PROMA modifies gradient accumulation across microbatches, the Intra-PROMA method defined below instead acts within each microbatch. Because the projection depends only on the current microbatch, Intra-PROMA is embarrassingly parallel across microbatches and compatible with standard data-parallel training.
1def proma_intra(token_advantages, act_in, grad_out, r=100, shrinkage=1.0):
2 """
3 Project out gradient components in the subspace spanned by
4 top-r singular vectors of activations and gradient outputs.
5 """
6 r = min(r, act_in.shape[0], act_in.shape[1], grad_out.shape[1])
7
8 Q_a = approx_rank_r_basis(act_in, r) # (d_in, r)
9 Q_g = approx_rank_r_basis(grad_out, r) # (d_out, r)
10
11 policy_grad = (token_advantages * grad_out.T) @ act_in / len(token_advantages)
12 projection = Q_g @ (Q_g.T @ policy_grad @ Q_a) @ Q_a.T
13 return policy_grad - shrinkage * projection
14
15def approx_rank_r_basis(X, r, power_iters=1):
16 """Approximate top-r right singular vectors via randomized SVD."""
17 n = X.shape[0]
18 omega = torch.randn(n, r)
19 Y = X.T @ omega
20 for _ in range(power_iters):
21 Y = X.T @ (X @ Y)
22 Q, _ = torch.linalg.qr(Y)
23 return Q
Figure 2: accumulate_PROMA is called during the backward pass of the policy gradient loss and modifies the gradient accumulation. Before adding the microbatch gradient, the partially accumulated gradient is projected to be orthogonal to the gradient of each sequence in the current microbatch (mcb_seq_grads).
To avoid the cost of operating on full gradient matrices, Intra-PROMA separately approximates the dominant subspaces of activations and gradient outputs. This is similar to the Kronecker factorizations which underlie K-FAC [8] and related second-order optimizers [12], which apply a sandwich product $G^{-1}(\cdot)A^{-1}$ to precondition the gradient. In contrast, PROMA uses the factored subspaces for projection rather than preconditioning. I have found that applying the sandwich projection to the subtracted component:
$\texttt{grad}-Q_{g}Q_{g}^{\top}\texttt{grad}\,Q_{a}Q_{a}^{\top}$ (1)
controls KL without impeding training progress. Alternatively one could apply a sandwich product of projections to the gradient (2) similarly to the $G^{-1}(\cdot)A^{-1}$ in K-FAC:
$(I-Q_{g}Q_{g}^{\top})\,\texttt{grad}\,(I-Q_{a}Q_{a}^{\top}),\qquad\neq\text{PROMA}$ (2)
but I found that (2) caused the validation score to plateau at a lower score.
4 Results
The demonstration was implemented in VeRL [11] (https://github.com/nilin/proma). Experiments used the GSM8K dataset [5] with the Qwen-3 0.6B model [13]. The PROMA configuration is based on REINFORCE without any PPO clipping of the likelihood ratio, to isolate the effect of the projected accumulation. All runs are without KL penalty. GRPO refers to the baseline with PPO clipping, and REINFORCE refers to the baseline without PPO clipping. All runs use the same group-relative advantage estimator. Figure 3 shows the performance and policy dynamics of PROMA compared to GRPO and REINFORCE.
(a) Validation performance
(b) KL vs. initial policy
(c) Entropy
(d) KL vs. lagged policy
Figure 3: Training performance and policy statistics for PROMA, GRPO, and REINFORCE. (a) Validation performance. (b) KL divergence from the initial policy. (c) Policy entropy. (d) KL divergence between the current policy and a lagged reference policy (running mean over 80 steps). Only GRPO (blue) uses PPO clipping of the likelihood ratio.
PROMA achieves comparable or improved validation performance relative to GRPO (Figure 3a). Although PROMA allows substantial deviation from the initial policy, its global KL divergence remains controlled compared to REINFORCE (Figure 3b). PROMA also maintains higher entropy for longer during training, indicating sustained exploration without premature policy collapse (Figure 3c). This is more pronounced with the accumulation-based PROMA rather than Intra-PROMA. Finally, PROMA has consistently lower KL divergence between successive policies than both baselines (Figure 3d), suggesting smoother local updates. These results indicate that PROMA achieves a favorable tradeoff between stability and flexibility, allowing meaningful policy improvement while enforcing local smoothness.
This experiment used the approximate complement projection shown in Algorithm 4. Additionally, the modification to the accumulated gradient was bounded: the subtracted component
$p=\texttt{t.grad}\allowbreak-\allowbreak\texttt{project\_to\_complement\allowbreak(t.grad)}$
was clamped to have norm at most $\|0.5*\texttt{mcb\_policy\_grad}\|$. This prevents the projection from removing more than a fixed fraction of the microbatch gradient’s magnitude.
1# Experiments used this approximate projection.
2def project_to_complement_v2(acc_vec, vecs):
3 U = vecs / torch.norm(vecs, dim = 0) # (d, k)
4 for _ in range(2):
5 for u in U.T:
6 acc_vec -= torch.dot(acc_vec, u) * u
7 return acc_vec
Figure 4: The results in this section used the iterative projection project_to_complement_v2 which takes $O(kd)$ FLOPS. I did not test if this was faster in practice.
5 Discussion
PROMA is a followup to ISOPO [1] which likewise aimed to define a proximal gradient update without a reference policy. In constrast to ISOPO which explored pre-conditioning based on re-scaling, PROMA instead projects away noisy directions to achieve proximal policy updates. The usefulness of projecting the accumulated gradient is consistent with findings in computational physics in which the SPRING optimizer [6] applied a projected momentum scheme to the MinSR [4] optimizer to achieve increased accuracy. Both MinSR and SPRING are variants of the variational Monte Carlo (VMC) method [2] for molecular energy calculation, a method which is formally similar to policy optimization. This parallel highlights a shared geometric structure underlying stochastic optimization methods in both computational physics and reinforcement learning.
6 Acknowledgements
I thank Molei Tao, Michael Luo, Gil Goldshlager, Jiahao Yao, Lin Lin, and Iulian Serban for inspiring conversations and encouragement.