### v4

# PROMA: Projected Microbatch Accumulation for Reference-Free Proximal Policy Updates

Abstract
This note introduces Projected Microbatch Accumulation (PROMA), a reference-free proximal policy method that controls KL divergence by projecting away high-variance components of the policy gradient. Two variants are presented. In the accumulation-based variant, the running gradient is projected orthogonal to the sequence-wise log-probability gradients of each microbatch. In the intra-microbatch variant, a factored projection using dominant subspaces of activations and gradient outputs is applied independently within each microbatch, making it compatible with standard data-parallel training. Empirically, the accumulation variant achieves tighter per-step KL control than GRPO with PPO clipping, while the intra-microbatch variant achieves the best validation performance.
1 Background
Proximal Policy Optimization (PPO) [10] and its variants stabilize reinforcement learning by clipping the policy likelihood ratio to limit deviations from a reference policy $\pi_{\text{old}}$. GRPO [11, 15] combines PPO-style clipping with a group-relative advantage estimator and has become the dominant approach for LLM fine-tuning. CISPO [4] is another variant that clips the importance-sampling ratio while ensuring that all tokens contribute to the gradient. More broadly, proximal policy methods can be viewed as tractable approximations to the natural policy gradient [8], whose exact computation is infeasible at scale.
2 Accumulation-based PROMA
Accumulation-based PROMA modifies the gradient during the backward pass. For each microbatch, the sequence-wise log-probability gradients (mcb_seq_grads in Algorithm 1) are computed as part of evaluating the policy gradient, denoted mcb_policy_grad. Before gradients are accumulated, the running gradient t.grad of a parameter tensor t is updated by projecting out components that lie in the span of the sequence-wise log-probability gradients of the current microbatch.
1def PROMA_accumulate(t, mcb_policy_grad, mcb_seq_grads)
2 """
3 k: number of sequences in microbatch
4 t: (d,) tensor of model weights in the layer
5 mcb_policy_grad: (d,) vanilla policy gradient w.r.t. tensor t for the microbatch
6 mcb_seq_grads: (d, k), sequence-wise gradients of log_prob.sum() w.r.t tensor t
7 """
8 complemented_acc_grad = project_to_complement(t.grad, mcb_seq_grads)
9 t.grad = complemented_acc_grad + mcb_policy_grad
10
11def project_to_complement(acc_vec, vecs):
12 Q, _ = torch.linalg.qr(vecs, mode="reduced") # Q: (d, k)
13 return acc_vec - Q @ (Q.T @ acc_vec)
Algorithm 1 PROMA_accumulate is called during the backward pass of the policy gradient loss and modifies the gradient accumulation. Before adding the microbatch gradient, the partially accumulated gradient is projected to be orthogonal to the gradient of each sequence in the current microbatch (mcb_seq_grads).
The parameter update $d\theta$ incurs a KL divergence that grows with the squared overlaps $\mathbb{E}_{s}[(d\theta\;\cdot\;\nabla_{\theta}\log\pi(s))^{2}]$, where $\nabla\log\pi(s)$ is the direction of fastest probability change for sequence $s$. PROMA controls this by projecting away these components of the gradient update.
Efficiency
To exactly project away the $k$ sequences in the microbatch takes $\approx 2k^{2}d$ FLOPs where $d$ is the number of model parameters. If the microbatch size is large then this can be applied group-wise so that $k\sim 8$. The cost $2k^{2}d$ of PROMA is then negligible compared to the $T\times d$ operations required to form the gradient, where $T$ is the number of tokens. Another option is to apply an approximate iterative projection which takes $O(kd)$ operations.
3 Intra-microbatch PROMA
While PROMA modifies gradient accumulation across microbatches, the Intra-PROMA method defined below instead acts within each microbatch. Because the projection depends only on the current microbatch, Intra-PROMA is embarrassingly parallel across microbatches and compatible with standard data-parallel training.
1def proma_intra(token_advantages, act_in, grad_out, r=100, shrinkage=1.0):
2 """
3 Project out gradient components in the subspace spanned by
4 top-r singular vectors of activations and gradient outputs.
5 """
6 r = min(r, act_in.shape[0], act_in.shape[1], grad_out.shape[1])
7
8 Q_a = approx_rank_r_basis(act_in, r) # (d_in, r)
9 Q_g = approx_rank_r_basis(grad_out, r) # (d_out, r)
10
11 policy_grad = (token_advantages * grad_out.T) @ act_in / len(token_advantages)
12 projection = Q_g @ (Q_g.T @ policy_grad @ Q_a) @ Q_a.T
13 return policy_grad - shrinkage * projection
14
15def approx_rank_r_basis(X, r, power_iters=1):
16 """Approximate top-r right singular vectors via randomized SVD."""
17 n = X.shape[0]
18 omega = torch.randn(n, r)
19 Y = X.T @ omega
20 for _ in range(power_iters):
21 Y = X.T @ (X @ Y)
22 Q, _ = torch.linalg.qr(Y)
23 return Q
Algorithm 2 proma_intra computes the Intra-PROMA gradient for a single layer within a microbatch. The projection subtracts gradient components lying in the dominant subspaces of activations and gradient outputs, approximated via randomized SVD.
The factored subspace approximation is similar to the Kronecker factorizations underlying K-FAC [9] and related second-order optimizers [13], which apply a sandwich product $G^{-1}(\cdot)A^{-1}$ to precondition the gradient. In contrast to K-FAC, PROMA applies a sandwich projection to the subtracted component:
$\texttt{grad}-Q_{g}Q_{g}^{\top}\texttt{grad}\,Q_{a}Q_{a}^{\top},$ (1)
meaning that intra-PROMA removes the intersection of the leading row and column substraces to control KL without impeding training progress. Alternatively one could apply a sandwich product of projections to the gradient (2) similarly to the $G^{-1}(\cdot)A^{-1}$ in K-FAC:
$(I-Q_{g}Q_{g}^{\top})\,\texttt{grad}\,(I-Q_{a}Q_{a}^{\top}).$ (2)
(2) would correspond to removing the union of leading row and column subspaces, and I found that this hurt validation performance compared to (1).
4 Results
Experimental setup
The demonstration was implemented in VeRL [12] (https://github.com/nilin/proma). The model is trained on MBPP [3] and validated on HumanEval [6], two related but distinct code generation benchmarks, using the Qwen-3 0.6B model [14]. The training batch consists of 64 prompts, each generating $n=16$ rollout samples. Policy updates use a mini-batch size of 32 and micro-batch size of 8, yielding $k=4$ microbatches per gradient accumulation step. All methods use a single PPO epoch per batch, the AdamW optimizer with weight decay $0.01$ and gradient clipping at $1.0$, and a constant learning rate schedule. Learning rates are swept over $\{2\times 10^{-6},\;5\times 10^{-6},\;1\times 10^{-5}\}$. The maximum prompt length is 512 tokens and the maximum response length is 1024 tokens. No KL penalty is applied in any configuration. Validation is performed every 2 training steps for 20 steps. All methods use the same GRPO group-relative advantage estimator [11].
Baselines and naming
This experiment compares four configurations. GRPO (w/ PPO-clip) uses PPO-style clipping of the likelihood ratio with clip ratio $\varepsilon=0.2$. GRPO (no PPO-clip) uses the same group-relative advantage but without clipping the likelihood ratio. PROMA (intra) applies Intra-PROMA (Algorithm 2) with projection dimension $d=10$. PROMA (accumulation) applies the accumulation-based PROMA (Algorithm 1) on the last 2 out of 4 microbatches. The PROMA configurations do not use PPO clipping, isolating the effect of the projected accumulation.
Two validation metrics are reported: the best validation score (peak score over the course of each run) and the final validation score (score at the last step), both aggregated as the median across runs.
(a) Best validation score vs. learning rate
(b) Final validation score vs. learning rate
(c) Best validation score per learning rate
(d) Final validation score per learning rate
Figure 1: MBPP $\to$ HumanEval validation results (Qwen-3 0.6B). (a–b) Learning rate sensitivity; each marker shows the median across seeds, $n$ indicates number of runs. (c–d) Per-learning-rate breakdown; each bar shows the median, $n$ denotes the number of seeds.
(a) Per-step KL divergence (smoothed)
(b) KL divergence from initial policy
Figure 2: KL metrics at lr $5\times 10^{-6}$. Lines show median, shaded regions show IQR.
Figure 1a–b shows learning rate sensitivity. PROMA (intra) achieves the highest best validation score at lr $5\times 10^{-6}$, and PROMA (accumulation) is the most robust to higher learning rates, maintaining strong performance at lr $1\times 10^{-5}$ where both GRPO variants degrade. For final validation scores, GRPO (no PPO-clip) drops sharply at lr $1\times 10^{-5}$, while GRPO (w/ PPO-clip) and PROMA (accumulation) remain more stable. Figure 1c–d provides a per-learning-rate breakdown.
Figure 2 shows KL dynamics at lr $5\times 10^{-6}$. GRPO (no PPO-clip) exhibits the largest per-step KL divergence (Figure 2a), while both PROMA variants maintain tighter local KL control, comparable to GRPO (w/ PPO-clip). PROMA (accumulation) achieves the lowest per-step KL overall. In terms of cumulative drift from the initial policy (Figure 2b), GRPO (w/ PPO-clip) stays closest to the initial policy, with PROMA variants in between, drifting less than GRPO (no PPO-clip) while still allowing more policy movement than the clipped baseline. PROMA (intra) achieves the best validation performance across all methods, while PROMA (accumulation) provides the tightest local KL control, both without relying on likelihood ratio clipping.
5 Discussion
PROMA is a followup to ISOPO [1] which likewise aimed to define a proximal gradient update without a reference policy. In contrast to ISOPO which explored pre-conditioning based on re-scaling, PROMA instead projects away noisy directions to achieve proximal policy updates. The usefulness of projecting the accumulated gradient is consistent with findings in computational physics in which the SPRING optimizer [7] applied a projected momentum scheme to the MinSR [5] optimizer to achieve increased accuracy. Both MinSR and SPRING are variants of the variational Monte Carlo (VMC) method [2] for molecular energy calculation, a method which is formally similar to policy optimization. This parallel highlights a shared geometric structure underlying stochastic optimization methods in both computational physics and reinforcement learning.
6 Acknowledgements
I thank Molei Tao and Michael Luo for inspiring conversations and encouragement.